# Boosting Generalization in Diffusion-Based Neural Combinatorial Solver via Energy-guided Sampling

### v1

@@ 4.2 Energy-guided Sampling for Problem Transfer
where $\mathcal{E}(y^{*}_{G^{\prime}},\mathbf{x}_{t};G^{\prime})=|y^{*}_{G^{\prime}}-f(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})|$ measures the energy between the optimal value and the predicted solution. Here, $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t})$ represents the predicted clean sample from the current noisy state $\mathbf{x}_{t}$. To overcome this issue, we parameterize the model outputs as the logits of $N$ independent Bernoulli samples, and estimate $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t})=\mathbb{E}_{\widetilde{\mathbf{x}}_{0}\sim p_{\theta}(\widetilde{\mathbf{x}}_{0}|\mathbf{x}_{t})}[\widetilde{\mathbf{x}}_{0}]$. Combining equations (10) and (11), we derive an energy-guided reverse sampling process:
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime})\propto p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G^{\prime})p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime}),$ (12)
$p(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})=\exp\left(-\frac{1}{\tau}\nabla_{\mathbf{x}_{t}}f\left(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}\right)\right).$
4.3 Analysis of Problem Transfer
@@
Consider the scenario where we aim to transfer knowledge from the original Traveling Salesman Problem (TSP) to its variants: the Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP). The diffusion model, trained on optimal TSP instances, estimates $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ such that, under the assumption of perfect optimization, $p_{\theta}(\mathbf{x}_{0}|G^{\prime})$ approximates $q_{\textup{TSP}}(\mathbf{x}_{0}|G^{\prime})$, where the latter represents the distribution of ground-truth TSP solutions for instance $G^{\prime}$. Through the Bayesian equation established in (10), we expect such a pre-trained prior would substantially enhance posterior sampling solution quality, as TSP shares fundamental structural similarities with PCTSP and OP, thereby encoding relevant domain knowledge. We formalize such similarities in the following analysis. The complete proof is provided in Appendix A.1. We make use of the following definition to facilitate the analysis.
@@
For a non-empty subset of nodes $S\subseteq V$, let $\textup{TSP}(S)$ and $\textup{argTSP}(S)$ denote the optimal cost and optimal tours of TSP on the subgraph specified by $S$. Under Assumptions A.2 and A.3, the optimal tours of PCTSP are $\textup{argTSP}(V\setminus S_{\textup{PCTSP}})$, where
@@
$S_{\textup{OP}}\in$ $\arg\min_{S\subseteq V}{\Delta(S)},\;\textup{ s.t. }\Delta(S)\geq\textup{TSP}(V)-D_{\textup{OP}}.$
@@
Theorem 4.2 shows that the optimal solutions of PCTSP and OP are indeed the optimal tours of TSP on some subgraph of $G^{\prime}$, revealing their fundamental structural similarities. We expect that the pre-trained diffusion model is powerful enough to generate high quality TSP solutions given test instances with various graph size. We can therefore understand the energy-guided sampling in this scenario as forcing the pre-trained model to focus on a subgraph of $G^{\prime}$. Through this theoretical analysis, we deepen our understanding on how the proposed training-free energy-guided sampling works when transferring from a pre-trained network to its variants.
5 Proposed Approach
Algorithm 1 Energy-guided Diffusion Sampling for Cross-problem Transfer
0:
1:  $p_{\theta}$: Pre-trained diffusion model
2:  $G^{\prime}$: Target problem instance
3:  $T$: Number of diffusion steps
4:  $\tau$: Energy guidance temperature
5:  $K$: Number of re-inference iterations
5:  Optimal solution $\mathbf{t}_{K}$ for problem instance $G^{\prime}$
6:  Initialize $\mathbf{x}_{T}$ with random binary values
7:  for $k=1$ to $K$ do
8:     Initialize $\mathbf{x}_{T}$ with previous best solution $\mathbf{t}_{k-1}$
9:     for $t=T$ to $1$ do
10:        Compute $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ from pre-trained model
11:        Compute energy gradient: $\nabla_{\mathbf{x}_{t}}f(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})$
12:        Compute $p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})\propto\exp(-\nabla_{\mathbf{x}_{t}}f/\tau)$
13:        Compute posterior: $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime})\propto p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G^{\prime})p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$
14:        Update next state with Bernoulli Sampling: $\mathbf{x}_{t-1}\sim\text{Cat}(\mathbf{x}_{t-1};p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime}))$
15:     end for
16:     Decode $\mathbf{x}_{0}$ to the best solution $\mathbf{t}_{k}$
17:  end for
18:  return  $\mathbf{t}_{K}$
Heatmap Generation. Building upon our theoretical analysis of the graph-based discrete diffusion model, we employ it to generate transition matrices for combinatorial optimization problems. To effectively process the input structure, we introduce an anisotropic graph neural network equipped with edge-gating mechanisms [37, 38]. This network encodes the input into a state matrix, where each element represents the selection probability $p_{\theta}(\mathbf{x}_{0}=1|G^{\prime})$ of nodes or edges as a heatmap representation. The pre-trained diffusion model then operates by progressively denoising a randomly perturbed graph structure to generate these probability heatmaps, which serve as the foundation for subsequent decoding steps.
Energy-guided Cross-problem Sampling. We propose a conditional energy-guided sampling framework that enhances cross-scale and cross-problem generalization without requiring additional training. We leverage DDIM [39] to accelerate the sampling process to 10 or 20 steps. While our initial experiments reveal that single-round sampling with energy function gradients yields suboptimal cross-problem performance, we address this limitation by adopting the rewrite technique from [7]. Specifically, we initialize each round by adding noise to the previous round’s best solution and iteratively apply energy-guided sampling (Algo. 1). This iterative process establishes a natural trade-off between solution quality and computational efficiency.
Decoding and Solution Selection. To construct feasible solutions from the generated heatmaps, we employ a greedy decoding strategy that iteratively selects nodes or edges with the highest probabilities until reaching termination conditions (e.g., forming a complete cycle in TSP applications). Our framework is also flexible in incorporating advanced decoding strategies to generate better results, such as 2-opt heuristics [40] and Monte Carlo Tree Search (MCTS) [22, 6].
Figure 2: Visualization of the node connectivity process. The first three images show the denoising steps of the pre-trained diffusion model for TSP solution generation, while the last shows the final PCTSP solution obtained through energy-guided sampling.
@@ 6 Numerical Results
Dataset. We evaluate our approach on the classical NP-complete combinatorial optimization problems: the Traveling Salesman Problem (TSP) together with its variants, the Prize Collecting Traveling Salesman Problem (PCTSP) and Orienteering Problem (OP).
@@
Evaluation. Following [9], we generate 1000 test instances for each problem scale: 20, 50, and 100, which denotes the node counts for PCTSP and OP. We evaluate model performance using two primary metrics: average solution cost and optimality gap relative to the exact solution. Additionally, we measure computational efficiency through total training time and per-instance inference time.
@@
Baselines. We compare our approach against multiple baseline categories. For PCTSP, (1) Exact solver: Gurobi; (2) OR-based heuristics: OR-Tools and Iterated Local Search (ILS); (3) Learning-based methods: AM [9], MDAM [41], AM-FT [11], ASP [10]. For OP, (1) Exact solver: Gurobi; (2) OR-based heuristics: Compass [42] and Tsili [43]; (3) Learning-based methods: AM [9], AM-FT [11].
@@
Our proposed DIF-Trans builds upon DIFUSCO’s TSP-trained checkpoints in three different scales without additional training. The results of DIF-Trans are recorded under 50 iterations with greedy decoding strategy. The inference steps are accelerated by DDIM [39] from 1000 training steps to 10 steps. The energy function used in guided sampling follows [35] which provides energy formulation for NP-complete CO problems, see A.2. The guidance temperature $\lambda$ is fixed to be 0.1 and the constraint coefficient $\beta$ is set to be 1.1. All experiments are conducted on a single Tesla V100 GPU.
6.1 Generalization Results
First, we conduct comprehensive generalization experiments following [10], examining our method’s transferability across both problem scales and variants. Our evaluation framework specifically examines: cross-scale generalization, where we assess performance across varying problem scales, and cross-problem generalization, focusing on adaptation to different problem variants. For cross-scale evaluation, we utilize a single model that exhibits optimal performance across all scales, eliminating the need for scale-specific training. Notably, while existing learning-based approaches require problem-specific retraining for PCTSP and OP instances, our DIF-Trans framework operates in a zero-shot generation.

### v2

@@ 4.1 Energy-guided Sampling for Problem Transfer
where $\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}))$ measures the distance of energy states between the optimal value and the predicted solution. Here, $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t})$ represents the predicted clean sample from the current noisy state $\mathbf{x}_{t}$, which is estimated by the logits of $N$ independent Bernoulli samples $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t})=\mathbb{E}_{\widetilde{\mathbf{x}}_{0}\sim p_{\theta}(\widetilde{\mathbf{x}}_{0}|\mathbf{x}_{t})}[\widetilde{\mathbf{x}}_{0}]$. In this work, we define $\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})):=\|y^{*}_{G^{\prime}}-\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})\|_{2}$. From Eq (12), we have:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ $\propto-\nabla_{\mathbf{x}_{t}}\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}),$ (13)
$p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ $=\exp\left(-\frac{1}{\tau}\nabla_{\mathbf{x}_{t}}\phi\left(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}\right)\right).$
where $\tau$ is the energy guidance temperature to control the strength of transferring. Combining Eq. (11) and Eq. (13), we derive an energy-guided reverse sampling process for problem transfer:
$\underbrace{p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime})}_{\text{Transfer Posterior}}\propto\underbrace{p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G^{\prime})}_{\text{Pre-trained Prior}}\underbrace{p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})}_{\text{Guidance Likelihood}}.$ (14)
In this work, we have the following log-barrier formulations for PCTSP and OP as the guided gradient:
Proposition 4.1 (Prize Collecting TSP (PCTSP)).
$\phi(\mathbf{x},\mathbf{y};G)$ $=f_{\text{cost}}(\mathbf{x},\mathbf{y};G)-\mu\cdot\log(-g(\mathbf{x},\mathbf{y};G))$
$g(\mathbf{x},\mathbf{y};G)$ $=\max(0,R-\sum_{v\in V}r_{v}y_{v})$
Proposition 4.2 (Orienteering Problem (OP)).
$\phi(\mathbf{x},\mathbf{y};G)$ $=f_{\text{cost}}(\mathbf{x},\mathbf{y};G)-\mu\cdot\log(-g(\mathbf{x},\mathbf{y};G))$
$g(\mathbf{x},\mathbf{y};G)$ $=\max(0,\sum_{e\in E}w_{e}x_{e}-B)$
4.2 Recursive Denoising-decoding Travel Strategy
Our initial experiments applying energy-guided denoising for cross-problem transfer highlighted a significant challenge: this approach alone often showed insufficient to effectively bridge the gap between the learned distribution of pre-trained problem types and the target distribution of new problem instances. This distributional divergence frequently hindered the generation of high-quality, feasible solutions for cross-problem transfer.
To address this, we first scaled up the inference time and explored a iterative rewrite strategy, adapting from [7]. This involved taking an initial solution $x_{0}^{(1)}$ for the pre-trained problem instance $G$, adding noise to obtain a highly noisy graph $x_{T}^{(1)}$, and then denoising to obtain the clean state $x_{0}^{(2)}$, aiming to progressively shift the solution towards the characteristics of the target problem. While this recursive strategy demonstrated improved transfer capabilities, it came at the cost of a considerable increase in inference time due to the repeated denoising iterations.
This led us to investigate the efficiency of the guidance mechanism itself. Inspired by the time-travel strategies in vision tasks [15], which revealed that guidance effects are not uniform across diffusion time steps, we hypothesized a similar phenomenon in NCO task. Our empirical analysis confirmed this: the impact of energy guidance varied significantly with the noise level. Crucially, we observed that guidance was most potent and beneficial at intermediate noise levels from 200 to 500 steps. In contrast, its effect was minimal on highly noisy graphs (analogous to an early time stage) and became largely redundant for nearly converged solutions (a late, refinement stage), where the solution structure is already well-defined.
Based on this observation that the impact of energy guidance at different noise levels, we developed our optimized recursive denoising-decoding travel strategy in Figure 3. Instead of applying full re-noising and iterative energy-guided denoising across many steps, our refined approach reduces the re-noising process to only a few steps. More importantly, the computationally intensive energy-guided denoising is applied for just a single step, precisely targeted at these effective intermediate noise levels. This targeted acceleration reduces the computational overhead associated with the recursive process, yielding a 5-10x speedup in inference time compared to the original iterations. This optimization not only maintains but enhances solution quality in several instances, establishing a significantly more favorable performance-to-speed trade-off.
Algorithm 1 outlines our proposed 2-phase Inference Time Adaptation framework, which combines recursive denoising-decoding travel strategy with energy-guided sampling. This approach offers a practical and efficient method for achieving zero-shot cross-problem transfer using pre-trained diffusion models, without any additional training costs. Furthermore, the framework exhibits flexibility by incorporating advanced heuristic decoding strategies, such as 2-opt heuristics [39] or Monte Carlo Tree Search (MCTS) [24, 6], to enhance solution quality.
Algorithm 1 Inference Time adaptation for Cross-problem Transfer
0:
1:  $p_{\theta}$: Pre-trained diffusion model
2:  $G^{\prime}$: Target problem instance
3:  $T$: Number of diffusion steps
4:  $i,j$: Number of intermediate steps
5:  $\tau$: Energy guidance temperature
6:  $K$: Number of recursive travel
6:  Optimal solution $\mathbf{y}_{K}$ for problem instance $G^{\prime}$
7:  Initialize $\mathbf{x}_{T}$ with random binary values
8:  for $k=1$ to $K$ do
9:     Re-noise previous estimated solution $\mathbf{y}_{k-1}$ to $\mathbf{x}_{j}$
10:     for $t=j$ to $i$ do
11:        Compute $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ from pre-trained model
12:        Compute energy gradient: $\nabla_{\mathbf{x}_{t}}\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})$
13:        Compute $p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})\propto\exp(-\nabla_{\mathbf{x}_{t}}\phi/\tau)$
14:        Compute posterior $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime})$ by Eq (14)
15:        Update next state with Bernoulli Sampling: $\mathbf{x}_{t-1}\sim\text{Cat}(\mathbf{x}_{t-1};p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime}))$
16:     end for
17:     Estimate $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{i})=\mathbb{E}_{\widetilde{\mathbf{x}}_{0}\sim p_{\theta}(\widetilde{\mathbf{x}}_{0}|\mathbf{x}_{i})}[\widetilde{\mathbf{x}}_{0}]$
18:     Decode $\widetilde{\mathbf{x}}_{0}$ to the optimal solution $\mathbf{y}_{k}$
19:  end for
20:  return  $\mathbf{y}_{K}$
4.3 Theoretical Analysis of Cross-problem Transfer
@@
Consider the scenario where we aim to transfer knowledge from the original Traveling Salesman Problem (TSP) to its variants: the Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP). The diffusion model, trained on optimal TSP instances, estimates $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ such that, under the assumption of perfect optimization, $p_{\theta}(\mathbf{x}_{0}|G^{\prime})$ approximates $q_{\textup{TSP}}(\mathbf{x}_{0}|G^{\prime})$, where the latter represents the distribution of ground-truth TSP solutions for instance $G^{\prime}$. Through the Bayesian equation established in (11), we expect such a pre-trained prior would substantially enhance posterior sampling solution quality, as TSP shares fundamental structural similarities with PCTSP and OP, thereby encoding relevant domain knowledge. We formalize such similarities in the following analysis. The complete proof is provided in Appendix B. We make use of the following definition to facilitate the analysis.
@@
For a non-empty subset of nodes $S\subseteq V$, let $\textup{TSP}(S)$ and $\textup{argTSP}(S)$ denote the optimal cost and optimal tours of TSP on the subgraph specified by $S$. Under Assumptions [^1] We assume simpler setups for PCTSP and OP for clearer presentation of the underlying similarities. These assumptions can be relaxed as discussed in Appendix C. B.2 and B.3, the optimal tours of PCTSP are $\textup{argTSP}(V\setminus S_{\textup{PCTSP}})$, where
@@
$S_{\textup{OP}}\in$ $\arg\min_{S\subseteq V}{\Delta(S)}$
[…]
@@
Theorem 4.4 shows that the optimal solutions of PCTSP and OP are indeed the optimal tours of TSP on some subgraph of $G^{\prime}$, revealing their fundamental structural similarities. We expect that the pre-trained diffusion model is powerful enough to generate high-quality TSP solutions given test instances with various graph sizes. We can therefore understand the inference time adaptation in this scenario as forcing the pre-trained model to focus on a subgraph of $G^{\prime}$. Through this theoretical analysis, we deepen our understanding of how the proposed training-free inference time adaptation works when transferring from a pre-trained network to its variants. Although we focus on PCTSP and OP in this paper, we show that our methodology and insight also hold for more challenging TSP variants such as TSP with time windows in Appendix D.
@@ 5 Numerical Results
Dataset. We evaluate our approach on the commonly-used NP-complete combinatorial optimization problems: the Traveling Salesman Problem (TSP) together with its variants, the Prize Collecting Traveling Salesman Problem (PCTSP) and Orienteering Problem (OP). The detailed problem descriptions are given in Appendix A.
@@
Evaluation. Following [10], we generate 1000 test instances for each problem scale: 20, 50, and 100, which denotes the node counts for PCTSP and OP. We evaluate model performance using two primary metrics: average solution cost and optimality gap relative to the exact solution solvers. Additionally, we measure computational efficiency through total training time and per-instance inference time.
@@
Baselines. We compare our approach against multiple baseline categories. For PCTSP, (1) Exact solver: Gurobi; (2) OR-based heuristics: OR-Tools and Iterated Local Search (ILS); (3) Learning-based methods: AM [10], MDAM [40], AM-FT [12], ASP [11], DIFUSCO [6], T2T [7]. For OP, (1) Exact solver: Gurobi; (2) OR-based heuristics: Compass [41] and Tsili [42]; (3) Learning-based methods: AM [10], AM-FT [12], DIFUSCO [6], T2T [7].
@@
Experimental Setup. The proposed Inference Time Adaptation for Diffusion-based solvers (DIFU-Ada) builds upon DIFUSCO’s [6] TSP-trained checkpoints in three different scales without additional training. The backbone archetechture used in diffusion model is the same as DIFUSCO, a 12-layer Anisotropic GNN with a width of 256 hidden nodes. The results of DIFU-Ada are recorded under 100 iterations, with 5-step re-noising and 1-step guided denoising. The decoding strategy used in experiments is greedy decoding. The energy functions used in guided sampling are stated in Proposition 4.1 and 4.2. The guidance temperature $\lambda$ is fixed to be 0.1 and the constraint coefficient $\mu$ is set to be 1 for all problem instances. All experiments are conducted on a single Tesla V100 GPU.
PCTSP-20 PCTSP-50 PCTSP-100
[…]
DIFUSCO 3.78 19.21% 1.04s 5.20 15.97% 1.35s 7.85 32.31% 2.02s
T2T 3.64 14.82% 2.02s 4.90 9.38% 2.18s 7.23 21.92% 2.70s
DIFU-Ada (Ours) 3.30 4.20% 1.72s 4.63 3.57% 1.99s 6.51 9.61% 2.74s
OP-20 OP-50 OP-100
[…]
DIFUSCO 9.25 12.48% 1.51s 25.60 13.45% 1.88s 45.66 20.02% 2.04s
T2T 9.67 8.51% 2.70s 26.89 9.09% 3.04s 48.72 14.70% 3.45s
DIFU-Ada (Ours) 10.24 3.11% 2.09s 28.21 4.63% 2.35s 54.56 8.06% 2.91s
Table 1: Zero-shot cross-problem transfer performance comparison between TSP-trained DIFUSCO, T2T and DIFU-Ada (Ours) approach on PCTSP and OP instances. The results show solution cost, optimality gap, and the average inference time for one instance across three problem scales (20, 50, and 100 nodes).
5.1 Cross-problem and Cross-scale Generalization Results
To validate the effectiveness of our inference time adaptation framework in zero-shot cross-problem transfer scenarios, we first conduct experiments while maintaining consistent problem scales. As demonstrated in Table 1, our method achieves significant zero-shot performance improvements when applied to the pre-trained DIFUSCO [6] and T2T [7] model on TSP. These improvements are evident across all PCTSP and OP problem scales, in both enhanced solution quality and reduced optimality gaps compared to the exact OR solvers.
Following the evaluation strategy following [11], we conducted experiments of DIFU-Ada’s transferability across diverse combinatorial optimization (CO) problem types and scales. Our evaluation framework specifically assesses two aspects of generalization: 1) Cross-problem Generalization: We assess the model’s adaptability to different CO problem variants, such as transferring from TSP to PCTSP or OP. 2) Cross-scale Generalization: We evaluate the performance of a single, optimally trained model across various problem scales (e.g., from small to large instances of the same problem type). Unlike existing learning-based methods that necessitate problem-specific retraining for new problem types or scales (e.g., for PCTSP and OP instances), our DIFU-Ada framework achieves zero-shot generation at inference time. This rigorous out-of-distribution evaluation strategy presents a more challenging benchmark than conventional in-distribution assessments, testing the model’s capacity of generalization.
The experimental results presented in Table 2 (PCTSP) and Table 3 (OP) show the cross-problem and cross-scale generalization capabilities of DIFU-Ada. Our framework achieves competitive average optimality gaps ($9.70\%$ for PCTSP and $7.26\%$ for OP) across all scales, while maintaining its zero-shot cross-problem transfer capability. This performance is achieved without the training costs (e.g., 3-5 days of retraining reported for existing methods) or the need for well-labeled training datasets. While OR-based heuristics like Compass exhibit strong performance on narrowly defined problem types, DIFU-Ada’s ability to leverage pre-trained diffusion models across different CO problem variants without additional training highlights its flexibility and broader applicability. Besides, the results on large-scale PCTSP-500 and PCTSP-1000 in Appendix F, Table 4 demonstrate the scalability of DIFU-Ada.